OpenStudy (anonymous):

Find the 10th term in the addition pattern with the formula: 8n – 3 A. 83 B. 77 C. 20 D. 15

OpenStudy (anonymous):

I think all you have to do is plug in 10 for n and solve.

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

Put 10 where n is so... \[8n-3\] becomes \[8(10) - 3 = 80 -3 = 77\]

OpenStudy (anonymous):

thank you